OpenStudy (anonymous):

If 3x^2-2x+7=0, then (x-1/3)^2=

OpenStudy (anonymous):

3x² - 2x + 7 = 0 3x² - 2x = - 7 x² - 2/3x = - 7/3 x² - 1/3x = - 7/3 + (1/3)² x² - 1/3x = - 21/9 + 1/9 Answer: (x - 1/3)² = - 20/9

OpenStudy (anonymous):

How did you get x^2-1/3x=-7/3+(1/3)^2?

OpenStudy (anonymous):

Thank you